How do you write the slope-intercept form of the equation of the line.
Through: (-3,-3), perpendicular to y= -3x

Answer by ewatrrr(24785)

 
Hi,
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
line.
Through: (-3,-3), perpendicular to y= -3x
New perpendicular Line m = 1/3 (negative reciprocal of -3)
y = (1/3)x + b |Using ordered pair (x,y):(-3-3) to solve for b
-3 = (1/3)*(-3) + b
-2 = b
y = (1/3)x -2